# PETSc/TAO Developments for Early Exascale Systems

### v1

@@ 1 Introduction
To program GPUs, developers can use vendor-provided programming models and libraries, such as CUDA, cuBLAS, and cuSparse from NVIDIA (NVIDA, 2024); HIP, hipBLAS, and hipSparse from AMD (AMD, 2024); and SYCL and OneAPI from Intel (Khronos SYCL Working Group, 2020). CUDA is proprietary and currently the most popular programming model for general-purpose processing on accelerators thanks to NVIDIA’s large market share. HIP, syntactically very similar to CUDA, allows developers to create portable applications for AMD and NVIDIA GPUs from single source code. On AMD GPUs, hipBLAS and hipSparse are simply wrappers over the rocBLAS and rocSparse libraries underneath from the AMD ROCm software stack for GPU programming. SYCL is a stark departure from CUDA and HIP and is advertised as an open standard allowing developers to use a C++ single source to program a wide range of targets, including CPUs, GPUs, digital signal processors (DSPs) and field programmable gate arrays (FPGAs). AMD and Intel also provide translation tools, such as hipify from AMD and SYCLomatic from Intel, to automatically convert source from CUDA to HIP or SYCL, respectively. Though these migration tools are helpful, none can guarantee automatic translation. Usually, programmers have to intervene to fix untranslatable code. To enhance performance portability across various GPUs and CPUs, ECP funded several projects such as RAJA (Beckingsale et al., 2019), Kokkos (Trott et al., 2022) and OpenMP target offload (OpenMP Architecture Review Board, 2021) compilers. RAJA and Kokkos use C++ templates and lambda functions to provide a portability layer in the front end but rely on vendor programming models and compilers in the backend. OpenMP target offload uses preprocessor directives (pragmas) in source code and thus requires compiler support. OpenMP has been a popular choice for Fortran applications, as Fortran support is poor in the other models. With all these options available, ECP applications and libraries were not required to use a certain programming model. The variety of programming languages, code complexity, varying package dependencies and large amount of legacy code precludes a one-size-fits-all solution.
@@
PETSc/TAO (PETSc for short) (Balay et al., 2024), one of the math libraries funded by ECP, features scalable solvers for nonlinear time-dependent differential and algebraic equations and for numerical optimization. PETSc is mainly written in the C programming language, though employing an object-oriented design. PETSc provides Fortran and Python bindings and interfaces with dozens of external packages in various languages, such as the direct solvers MUMPS (in Fortran) (Amestoy et al., 2001) and SuperLU (in C) (Demmel et al., 2024), the multigrid solver Hypre (in C) (Falgout, 2023), and the adaptive mesh refinement framework AMReX (in C++) (Zhang et al., 2019). PETSc, which has formed an ecosystem with applications, libraries and frameworks built upon it (Adams et al., 2022a), is widely used in academia, government laboratories, and industry and is an important building block for many simulation codes. Likewise, PETSc is part of the xSDK (xSDK Team, 2024) and E4S (E4S Team, 2024) ecosystems of complementary HPC packages.
@@
Because PETSc functions as numerical infrastructure for many application packages, the library design completely separates the programming models used by applications (or external packages) and the models used by PETSc for its backend computational kernels. This flexibility is accomplished by sharing data between the application and PETSc programming models, but not sharing the programming models’ internal data structures. Because the data is shared, there are no copies between models and no loss of efficiency. This separation allows PETSc users from C/C++, Fortran, or Python to employ their preferred GPU programming model, whether it be CUDA, HIP, or Kokkos. In all cases, users can rely on PETSc’s large assortment of composable, hierarchical, and nested solvers (Brown et al., 2012), as well as advanced time-stepping and adjoint capabilities (Zhang et al., 2022) and numerical optimization methods running on the GPU. For example, an application for solving time-dependent partial differential equations may compute the Jacobian using Kokkos and then call PETSc’s time-stepping routines and algebraic solvers that use CUDA. Applications can mix and match programming models, allowing, for example, some application code in HIP and some in CUDA.
@@
In Mills et al. (2021), we provided a blueprint for porting PETSc applications to use GPUs, surveyed challenges in developing efficient and portable mathematical libraries for GPU systems, and introduced the PETSc backend developments meant to meet these challenges at that time. With the sunset of ECP and the availability of early exascale computers such as Frontier and Aurora, we summarize the accomplishments of the PETSc/TAO ECP project, and we present the latest developments on performance portability through PETSc, illustrated by results on exascale computers. This paper is organized as follows. We first recap some GPU programming challenges and PETSc’s responses in Section 2, then provide an update in Section 3 to the PETSc application GPU-porting blueprint. Afterwards, we introduce in detail three low-level PETSc developments for GPUs, including communication, matrix assembly, and asynchronous solvers in Sections 4 to 6. Next, we move higher in the PETSc software hierarchy and present PETSc GPU developments for algebraic multigrid, batched linear solvers, and dense reformulations of the limited memory BFGS method (L-BFGS) (Liu and Nocedal, 1989) in Sections 7 to 9. In Section 10, we move to the top of software stack and present a mini-application case study, examining a fully GPU-enabled implementation of time evolution of the Landau collision integral using PETSc. We conclude in Section 11. Throughout the paper, we use the term programming model to refer to both the model and its supporting runtime. With one exception, our experiments were conducted on the four (pre-) exascale machines with configurations shown in Table 1.
@@ 2 GPU programming challenges and PETSc’s responses
    Algorithms for high-throughput systems: design parallel algorithms exploiting high GPU-concurrentcy.
@@
In addition, we talked about seven ancillary challenges:
@@
Roughly speaking, while F1 concerns portability, A5 and A6 are about correctness, and F2, F3, A1–A4 and A7 are related to performance. To meet these challenges, during ECP we greatly improved PETSc’s GPU capability around these three aspects with new GPU backends and constructs, new GPU-friendly application programming interfaces (APIs), and new algorithms for high-throughput computation. One can refer to Mills et al. (2021) for categorized PETSc responses to these challenges. Here we further explain the issues of portability and correctness with respect to the latest developments in PETSc.
@@
Before ECP, we had a legacy PETSc/CUDA backend for NVIDIA GPUs, and a PETSc/OpenCL backend via ViennaCL (Rupp et al., 2016) for OpenCL devices. We implemented vector and matrix sub-types for these backends, such as VECCUDA and MATAIJVIENNACL, and encouraged users to use PETSc’s options database from the command line, e.g., -vec_type cuda or -mat_type aijviennacl to set GPU-specific object types at runtime. Using these options, the same PETSc application source code can work with either a CUDA device or an OpenCL device. We did not write many device kernels ourselves. Instead we relied on vendor libraries, such as NVIDIA’s cuBLAS, cuSparse, etc. to provide basic vector and matrix operations at the process level. PETSc mainly managed the MPI parallelism and host-device synchronization. During ECP, we also needed to support AMD and Intel GPUs. Ideally we should use a unified, portable programming model for all devices, but unfortunately no suitable model existed at the time. OpenCL has portable performance as an overarching goal, but it is verbose and implementation quality lags behind the standard needed for many application codes (Pennycook et al., 2013). Therefore, for NVIDIA GPUs, we further optimized and expanded the PETSc/CUDA backend with new features. Later, staff from AMD helped us add a PETSc/HIP backend for AMD GPUs, which was very similar to the PETSc/CUDA backend thanks to the similarity between HIP and CUDA. We partially combined the PETSc/CUDA and the PETSc/HIP backends via a C++ abstraction layer we developed named CUPM (CUDA-like Programing Model). Further work is needed to fully combine the two backends. We did not use hipify, the AMD CUDA to HIP migration tool, because PETSc supports multiple versions of CUDA and contains many CUDA library calls to cuBLAS, cuSparse etc., which makes the translation more difficult. In addition to these vendor-native models, we developed a new GPU backend using the Kokkos programming model and employing Kokkos-Kernels (a math library providing BLAS, sparse BLAS etc.) for basic math operations. We currently do not have a RAJA implementation, as it lacks a math kernel library similar to Kokkos-Kernels. The PETSc/Kokkos backend is able to run on various CPUs and GPUs, including ones from NVIDIA, AMD and Intel. Going forward, as portable device programming models mature, we hope to consolidate these PETSc GPU backends to reduce maintenance burden and avoid code duplication. However, internal reorganization will not affect existing PETSc users working with PETSc objects. As mentioned before, we share data between an application and the PETSc programming model, but do not share the programming models’ internal data structures. We do provide utility APIs for PETSc users employing some popular GPU programming models. For example, for Kokkos users, we have APIs returning a Kokkos::View from a PETSc vector; for CUDA users, given the same PETSc vector, we have APIs returning a device pointer pointing to the vector storage in device memory.
@@
Streams are an important mechanism on GPUs to hide serial kernel launch latency by pipelining kernel launches. Streams can also improve resource utilization with extra parallelism by potentially executing kernels in different streams concurrently. In the latter sense, GPU streams are like CPU threads, but the programming paradigm is very different. Suppose that there are two functions on the host, a caller and a callee. With CPU threads, the callee is automatically executed on the same thread as caller. With GPU streams, the caller must pass the callee the stream that it uses by some mechanism, e.g., via an argument, through a global variable, or they may both assume a default stream. For an existing library written in C++, one might have the luxury to overload all its API functions with an extra stream argument without breaking existing code. Though doable, it is not only daunting but also error-prone because users might call an original API when they are supposed to call the one with a stream argument, resulting in mismatched streams being used. For a C library like PETSc, we cannotafford to break the API or bloat it with a host of new functions merely adding an additional stream argument. Thus, at the very basic level, we use a variable storing a PETSc global stream and wrap it into a generic object of type PetscDeviceContext. PetscDeviceContext provides uniform APIs across different PETSc GPU backends such as CUDA, HIP and SYCL. Users can query the device type to get a handle to the stream, and then cast the handle to the specific stream type according to the device type, as shown in the code below. For Kokkos users, we provide a utility function PetscGetKokkosExecutionSpace() to directly return a Kokkos execution space instance with the stream.
@@
From the user’s point of view, PETSc works as if using one global GPU stream. PETSc users should query this stream with the above methods when they want to coordinate with PETSc streams. The type of stream can be changed globally with the command line option -root_device_context_stream_type. We are exploring multi-stream parallelism with the PetscDeviceContext construct, see Section 6.2.
@@ 3 Porting PETSc applications to GPUs
This pattern holds whether targeting CPUs or GPUs for execution. When porting GPUs, the creation and manipulation of solver, matrix, and vector objects does not change, but users will need to 1) write code to ensure that needed data structures are either copied from CPU memory to device memory or constructed directly on the GPU and 2) provide Function and (optionally) Jacobian routines that call GPU kernels. We recommend adopting an incremental approach, in which portions of the computation are moved to GPU and then evaluated for correctness and performance; porting can be considered complete when observed GPU speedup is sufficient, relative to the cost of the rest of the simulation run.
@@
Listing 2 shows a Kokkos implementation of Function, which is similar to the traditional CPU version. For simplicity, we assume periodic boundary conditions in one dimension, but the pattern is similar in more dimensions and with general boundary conditions. DMDAVecGetKokkosOffsetView(da, xl, &X) returns a Kokkos OffsetView X from a PETSc vector xl. Because X and xl share data, we wrap but do not copy xl’s data. Kokkos OffsetViews are essentially multi-dimensional arrays with non-zero start indices. PETSc uses them so that users can access a locally owned array with conceptually simpler global indices. Via the mesh management object da, PETSc knows the dimension, start and end indices of the OffsetView so that the returned object is properly populated. Then, users can write parallel device code of their choice with these OffsetViews. A critical issue here is that users must be aware of the asynchronous nature of GPU computations, and therefore must know the current GPU stream used by PETSc. PETSc Kokkos users have the PetscGetKokkosExecutionSpace() to get a Kokkos execution space instance that wraps the stream. With that, one can construct a Kokkos RangePolicy object, for example in a Kokkos parallel_for dispatch, so that the user’s device code and PETSc’s share the same stream, and to enforce data dependency.
@@
The Jacobian computation departs from the block-oriented MatSetValues() approach traditionally used in CPU-based PETSc codes, as it is not possible to efficiently implement on GPUs. We discuss this topic in detail in Section 5, and here simply note that to use the new coordinate-based MatSetValuesCOO() approach each GPU thread places the non-zeros it contributes into a device buffer and later calls MatSetValuesCOO() to have PETSc post-process the buffer, as shown in Listing 3.

### v2

# PETSc/TAO Developments for GPU-Based Early Exascale Systems

@@ 1 Introduction
To program GPUs, developers can use vendor-provided programming models and libraries, such as CUDA, cuBLAS, and cuSparse from NVIDIA (NVIDA, 2024); HIP, hipBLAS, and hipSparse from AMD (AMD, 2024); and SYCL and OneAPI from Intel (Khronos SYCL Working Group, 2020). CUDA is proprietary and currently the most popular programming model for general-purpose processing on accelerators thanks to NVIDIA’s large market share. HIP, syntactically very similar to CUDA, allows developers to create portable applications for AMD and NVIDIA GPUs from a single source code. On AMD GPUs, hipBLAS and hipSparse are simply wrappers over the rocBLAS and rocSparse libraries underneath the AMD ROCm software stack for GPU programming. SYCL is a stark departure from CUDA and HIP and is advertised as an open standard allowing developers to use a C++ single source to program a wide range of targets, including CPUs, GPUs, digital signal processors (DSPs), and field programmable gate arrays (FPGAs). AMD and Intel also provide translation tools, such as hipify from AMD and SYCLomatic from Intel, to automatically convert source codes from CUDA to HIP or SYCL, respectively. Though these migration tools are helpful, none can guarantee automatic translation. Usually, programmers have to intervene to fix untranslatable code. To enhance performance portability across various GPUs and CPUs, ECP funded several projects such as RAJA (Beckingsale et al., 2019), Kokkos (Trott et al., 2022) and OpenMP target offload (OpenMP Architecture Review Board, 2021) compilers. RAJA and Kokkos use C++ templates and lambda functions to provide a portability layer in the front end but rely on vendor programming models and compilers in the backend. OpenMP target offload uses preprocessor directives (pragmas) in source code and thus requires compiler support. OpenMP has been a popular choice for Fortran applications, as Fortran support is poor in the other models. With all these options available, ECP applications and libraries were not required to use a certain programming model. The variety of programming languages, code complexity, varying package dependencies, and large amount of legacy code precludes a one-size-fits-all solution.
@@
PETSc/TAO (PETSc for short) (Balay et al., 2024), one of the math libraries funded by ECP, features scalable solvers for nonlinear time-dependent differential and algebraic equations and for numerical optimization. PETSc is mainly written in the C programming language, though employing an object-oriented design. PETSc provides Fortran and Python bindings and interfaces with dozens of external packages in various languages, such as the direct solvers MUMPS (in Fortran) (Amestoy et al., 2001) and SuperLU (in C) (Demmel et al., 2024), the multigrid solver Hypre (in C) (Falgout, 2023), and the adaptive mesh refinement framework AMReX (in C++) (Zhang et al., 2019). PETSc, which has formed an ecosystem with applications, libraries, and frameworks built upon it (Adams et al., 2022a), is widely used in academia, government laboratories, and industry and is an important building block for many simulation codes. Likewise, PETSc is part of the xSDK (xSDK Team, 2024) and E4S (E4S Team, 2024) ecosystems of complementary HPC packages.
@@
Because PETSc functions as numerical infrastructure for many application packages, the library design completely separates the programming models used by applications (or external packages) and the models used by PETSc for its backend computational kernels. This flexibility is accomplished by sharing data between the application and PETSc programming models but not sharing the programming models’ internal data structures. Because the data is shared, there are no copies between models and no loss of efficiency. This separation allows PETSc users from C/C++, Fortran, or Python to employ their preferred GPU programming model, whether it be CUDA, HIP, or Kokkos. In all cases, users can rely on PETSc’s large assortment of composable, hierarchical, and nested solvers (Brown et al., 2012), as well as advanced time-stepping and adjoint capabilities (Zhang et al., 2022) and numerical optimization methods running on the GPU. For example, an application for solving time-dependent partial differential equations may compute the Jacobian using Kokkos and then call PETSc’s time-stepping routines and algebraic solvers that use CUDA. Applications can mix and match programming models, allowing, for example, some application code in HIP and some in CUDA.
@@
In Mills et al. (2021), we provided a blueprint for porting PETSc applications to use GPUs, surveyed challenges in developing efficient and portable mathematical libraries for GPU systems, and introduced the PETSc backend developments meant to meet these challenges at that time. With the sunset of ECP and the availability of early exascale computers such as Frontier and Aurora, we summarize the accomplishments of the PETSc/TAO ECP project, and we present the latest developments on performance portability through PETSc, illustrated by results on exascale computers. This paper is organized as follows. We first recap some GPU programming challenges and PETSc’s responses in Section 2, then provide an update in Section 3 to the PETSc application GPU-porting blueprint. Afterwards, we introduce in detail some lower-level PETSc developments for GPUs, including improved communication support (including asynchronous, GPU-initiated communication) and efficient matrix assembly in Sections 4 and 5. Next, we move higher in the PETSc software hierarchy and discuss two approaches that can increase GPU throughput and utilization for linear solves: asynchronous solvers and batched solvers (Sections 6 and 7, respectively). We then discuss the use of GPU-friendly dense reformulations of algorithms to improve the performance of the limited memory BFGS method (L-BFGS) in Section 8. The paper culminates in two mini-application studies using PETSc on GPUs: solving the 3D Laplacian using algebraic multigrid (Section 9) and time evolution of the Landau collision integral (Section 10). We present some brief conclusions and thoughts towards the future in Section 11. We note that, throughout the paper, we use the term programming model to refer to both the model and its supporting runtime. With one exception, our experiments were conducted on the four (pre-) exascale machines with configurations shown in Table 1.
@@ 2 GPU programming challenges and PETSc’s responses
    Algorithms for high-throughput systems: design parallel algorithms exploiting high GPU-concurrency.
@@
In addition, we discussed seven ancillary challenges:
@@
While F1 concerns portability, A5 and A6 are about correctness, and F2, F3, A1–A4 and A7 are related to performance. To meet these challenges, during ECP we greatly improved PETSc’s GPU capability around these three aspects with new GPU backends and constructs, new GPU-friendly application programming interfaces (APIs), and new algorithms for high-throughput computation. One can refer to Mills et al. (2021) for categorized PETSc responses to these challenges. Here, we further explain the issues of portability, correctness, and multiprecision with respect to the latest developments in PETSc.
@@
Before ECP, we had a legacy PETSc/CUDA backend for NVIDIA GPUs, and a PETSc/OpenCL backend via ViennaCL (Rupp et al., 2016) for OpenCL devices. We implemented vector and matrix sub-types for these backends, such as VECCUDA and MATAIJVIENNACL, and encouraged users to use PETSc’s options database from the command line, e.g., -vec_type cuda or -mat_type aijviennacl to set GPU-specific object types at runtime. Using these options, the same PETSc application source code can work with either a CUDA device or an OpenCL device. We did not write many device kernels ourselves. Instead, we relied on vendor libraries to provide basic vector and matrix operations at the process level. PETSc mainly managed the MPI parallelism and host-device synchronization. During ECP, we also needed to support AMD and Intel GPUs. Ideally, we should use a unified, portable programming model for all devices, but unfortunately, no suitable model existed then. OpenCL has portable performance as an overarching goal, but the API is verbose, and vendor support for OpenCL in compilers and performance tools is limited (Bertoni et al., 2020). Therefore, for NVIDIA GPUs, we further optimized and expanded the PETSc/CUDA backend with new features. Later, staff from AMD helped us add a PETSc/HIP backend for AMD GPUs, which was very similar to the PETSc/CUDA backend thanks to the similarity between HIP and CUDA. We partially consolidated the PETSc/CUDA and the PETSc/HIP backends via a C++ abstraction layer we developed named CUPM (CUDA-like Programing Model) , which hides CUDA/HIP API disparities in PETSc. Further work is needed to fully consolidate the two backends. We did not use hipify, the AMD CUDA to HIP migration tool, because PETSc supports multiple versions of CUDA and contains many CUDA library calls to cuBLAS and cuSparse, which makes the translation more difficult. In addition to these vendor-native models, we developed a new GPU backend using the Kokkos programming model and employing Kokkos-Kernels (a math library providing BLAS, sparse BLAS etc.) for basic math operations. We currently do not have a RAJA implementation, as it lacks a math kernel library similar to Kokkos-Kernels. The PETSc/Kokkos backend can run on various CPUs and GPUs, including NVIDIA, AMD, and Intel.
Going forward, as portable device programming models mature, we hope to consolidate these PETSc GPU backends to reduce maintenance burden and avoid code duplication. However, internal reorganization will not affect existing PETSc users working with PETSc objects. As mentioned before, we share data between an application and the PETSc programming model, but do not share the programming models’ internal data structures. We do provide utility APIs for PETSc users employing some popular GPU programming models. For example, for Kokkos users, we have APIs returning a Kokkos::View from a PETSc vector; for CUDA users, given the same PETSc vector, we have APIs returning a device pointer pointing to the vector storage in device memory.
@@
Streams are an important mechanism on GPUs to hide serial kernel launch latency by pipelining kernel launches. Streams can also improve resource utilization with extra parallelism by potentially executing kernels in different streams concurrently. In the latter sense, GPU streams are like CPU threads, but the programming paradigm is very different. Suppose that there are two functions on the host, a caller and a callee. With CPU threads, the callee is automatically executed on the same thread as the caller. With GPU streams, the caller must pass the callee the stream that it uses by some mechanism, e.g., via an argument, through a global variable, or they may both assume a default stream. For an existing library written in C++, one might have the luxury of overloading all its API functions with an extra stream argument without breaking existing code. Though doable, it is not only daunting but also error-prone because users might call an original API when they are supposed to call the one with a stream argument, resulting in mismatched streams being used. For a C library like PETSc, we cannot afford to break the API or bloat it with a host of new functions merely by adding an additional stream argument. Thus, at the very basic level, we use a variable storing a PETSc global stream and wrap it into a generic object of type PetscDeviceContext. PetscDeviceContext provides uniform APIs across different PETSc GPU backends such as CUDA, HIP and SYCL. Users can query the device type to get a handle to the stream, and then cast the handle to the specific stream type according to the device type, as shown in the code below. For Kokkos users, we provide a utility function PetscGetKokkosExecutionSpace() to directly return a Kokkos execution space instance with the stream.
@@
From the user’s point of view, PETSc works as if using one global GPU stream. (If we find that users wish to make PETSc calls from multi-threaded GPU user code, we could add an option to protect this global state with locks or to store it in thread-local storage, analogously to how an OpenMP thread-safe subset of PETSc can be configured with the --with-threadsafety option). PETSc users should query this stream with the above methods when they want to coordinate with PETSc streams. The type of stream can be changed globally with the command line option -root_device_context_stream_type. We are exploring multi-stream parallelism with PETSc device contexts, see Section 6.2.
On PETSc’s multiprecision support:
Partly because of PETSc’s origins as a C library, it currently supports only one precision that is set at configuration time. This precision—which could be single, double, float128, or fp16—defines PetscScalar, the data type of all PETSc vector and matrix entries. Enforcing a uniform data type simplifies the code design and helps to prevent bugs and improve performance. Some users, however, need to call certain external packages (e.g., the direct solver SuperLU (Demmel et al., 2024)) in lower precision to save memory. Furthermore, the proliferation of GPUs supporting higher performance in lower precisions makes it desirable to enable some multiprecision support in PETSc. For a few external packages such as SuperLU and HPDDM (Jolivet et al., 2021), we have already provided options to use a precision other than PetscScalar. For example, one can use -pc_factor_mat_solver_type superlu_dist -pc_precision single to call SuperLU in single while PETSc is in double precision. PETSc will do the precision conversion automatically when building data structures for the packages. We can adopt the same approach when offloading data to GPUs. As mentioned above, PETSc maintains two copies of data: one on the host and the other on the device. Currently, they are in the same precision. We plan to provide new APIs to let users specify a different precision for data on device for PETSc GPU Vec/Mat types, either globally or at the linear solver level. This will allow some multiprecision support for GPUs, without greatly increasing code complexity and without affecting the CPU user API.
@@ 3 Porting PETSc applications to GPUs
This pattern holds whether targeting CPUs or GPUs for execution. When porting to GPUs, the creation and manipulation of solver, matrix, and vector objects do not change, but users will need to 1) write code to ensure that needed data structures are either copied from CPU memory to device memory or constructed directly on the GPU and 2) provide Function and (optionally) Jacobian routines that call GPU kernels. We recommend adopting an incremental approach, in which portions of the computation are moved to GPU and then evaluated for correctness and performance; porting can be considered complete when observed GPU speedup is sufficient, relative to the cost of the rest of the simulation run.
@@
Listing 2 shows a Kokkos implementation of Function, which is similar to the traditional CPU version. For simplicity, we assume periodic boundary conditions in one dimension, but the pattern is similar in more dimensions and with general boundary conditions. DMDAVecGetKokkosOffsetView(da, xl, &X) returns a Kokkos OffsetView X from a PETSc vector xl. Because X and xl share data, we wrap but do not copy xl’s data. Kokkos OffsetViews are essentially multi-dimensional arrays with non-zero start indices. PETSc uses them so that users can access a locally owned array with conceptually simpler global indices. Via the mesh management object da, PETSc knows the dimension, start, and end indices of the OffsetView so that the returned object is properly populated. Then, users can write parallel device code of their choice with these OffsetViews. A critical issue here is that users must be aware of the asynchronous nature of GPU computations, and therefore must know the current GPU stream used by PETSc. PETSc Kokkos users have the PetscGetKokkosExecutionSpace() to get a Kokkos execution space instance that wraps the stream. With that, one can construct a Kokkos RangePolicy object, for example in a Kokkos parallel_for dispatch, so that the user’s device code and PETSc’s share the same stream, and to enforce data dependency.
@@
The Jacobian computation departs from the block-oriented MatSetValues() approach traditionally used in CPU-based PETSc codes, as it does not map efficiently onto GPU architectures. We discuss this topic in detail in Section 5, and here simply note that to use the new coordinate-based MatSetValuesCOO() approach each GPU thread places the non-zeros it contributes into a device buffer and later calls MatSetValuesCOO() to have PETSc post-process the buffer, as shown in Listing 3.